Find x. e^π = e^(2x) sinx


1
Expert's answer
2020-12-24T11:09:54-0500

Here given that ,"e^\pi=e^{2x}.sinx"

"\implies e^{\pi-2x}=sinx" ......(1)

Again we know that the value of "sinx" lies between "(-1)" and "1" .

Therefore, "-1\leq sinx\leq1"

"\implies-1\leq e^{\pi-2x}\leq1" [from (1)]


Because "-1\leq e^{\pi-2x}" is always true, the case "e^{π-2x} \leq 1" will be considered.

"e^{\pi-2x}\leq1"

"\implies" "e^{\pi-2x}\leq e^0"

"\implies(\pi-2x)\leq 0" [As "e^{\pi-2x}" is a decreasing function]

"\implies x\geq \frac{\pi}{2}" .......(2)

Because "e^{\pi-2x}\geq 0" , "\sin(x) \geq0," then "0 \leq x \leq \pi."

But here "x<\frac{\pi}{2}" does not satisfy the inequality (2).

Therefore "x=\frac{\pi}{2}" is the required solution of the given equation.